how many moles of calcium is required to write April in chalk
Answer #1

Actually, if you're given just that, then you assume that it is one mole of the compound. If you want to know how many grams there are in one mole of that compound, you add up the molar mass of each element,
Ca=40.08
C=12.01
0= 3(16.00) because there are three oxygen atoms
so, you get 40.08+48.00+12.01
= 88.08+12.01
= 100.09g/mole
If you want the number of particles you multiply by Avogadro's number (6.022x10^(23) particles/mole)

to calculate the moles you need the initial weight of chalk and the weight after writing april. minus both the weights and now use this weight W=w1-w2 for calculating moles.

for explaning- assume you have 150 grams of CaCO3, you can calculate how many moles of CaCO3 you have by dividing the mass of CaCO3 by the molar mass of CaCO3, like this...

150 grams / 100 grams/mole = 1.5 moles

I hope that you like this answer. If so, please vote for it as the "Best Answer".